OpenStudy (anonymous):

Factor completely: 3x2 + 2x − 1 (3x + 1)(x − 1) (3x + 1)(x + 1) (3x − 1)(x + 1) (3x − 1)(x − 1)

OpenStudy (nurali):

x² + 1/3x = 1/3 + (1/3)² x² + 1/3x = (3 + 1)/9 (x + 1/3)² = 4/9 x + 1/3 = ± 2//3 = x + 1/3 - 2/3, = x - 1/3, = 3x - 1 = x + 1/3 + 2/3, = x + 3/3, = x + 1